OpenStudy (anonymous):

The volume of a cone is 183.26 cubic feet and its radius is 5 feet. What is the cone's height? Round to the nearest foot. Use 3.14 for \pi .

OpenStudy (anonymous):

@DaBest21

OpenStudy (michele_laino):

here we have to use this formula: \[\Large V = \frac{{\pi {r^2}h}}{3}\]

OpenStudy (michele_laino):

where V is the volume of the cone, r is the radius of its base, and h is the height of the cone

OpenStudy (anonymous):

\[3\frac{ 183.26}{ 3.14*5*5 }\]is the formula @Aliypop

OpenStudy (anonymous):

so do 3.14*5*5 first @Aliypop

OpenStudy (michele_laino):

from that formula above we get: \[h = \frac{{3V}}{{\pi {r^2}}} = \frac{{3 \times 183.26}}{{3.14 \times 5 \times 5}} = ...feet\]

OpenStudy (anonymous):

78.5

OpenStudy (anonymous):

yep now do 183.26/78.5

OpenStudy (anonymous):

2.33452293

OpenStudy (michele_laino):

hint: \[h = \frac{{3V}}{{\pi {r^2}}} = \frac{{3 \times 183.26}}{{3.14 \times 5 \times 5}} = 3 \times 2.335 = ...feet\]

OpenStudy (anonymous):

2.3335

OpenStudy (anonymous):

what's the next step ?

OpenStudy (michele_laino):

as I wrote before, what is: 3*2.335=...?

OpenStudy (anonymous):

7.005

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

so7.005 is the answer?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

no

OpenStudy (anonymous):

u have to round

OpenStudy (anonymous):

okay :D

OpenStudy (anonymous):

@Aliypop round

OpenStudy (michele_laino):

yes! sorry @Aliypop you have to round off to the hearest unit of foot

OpenStudy (michele_laino):

nearest*

OpenStudy (anonymous):

so what does 0 tell u to do to 7 Hint less than 5 means stay the same

OpenStudy (anonymous):

so would it be 8.00

OpenStudy (anonymous):

no it would be 7 Don't forget to medal and fan if you already haven't. Plus send me a message if you need more help @Aliypop

OpenStudy (anonymous):

Carson's Bakery and Chocolate Shop is experimenting with making chocolate in the shape of different types of sports equipment. Bayside High School's Volleyball team has ordered 100 solid white chocolate volleyballs. What volume of white chocolate is needed to fill this order if the radius of each volleyball is 4 cm? Rounded to the nearest tenth of a cubic centimeter. Use 3.14 for \pi .

OpenStudy (anonymous):

it appeared wrong again

OpenStudy (anonymous):

we did this already didnt we? @Aliypop

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@welshfella @welshfella

OpenStudy (anonymous):

I got 211 now as the answer

OpenStudy (michele_laino):

hint: the volume V of a sphere is given by the subsequent equation (pi=3.14): \[\begin{gathered} V = \frac{{4 \times 3.14}}{3}{r^3} = \frac{{12.56}}{3} \times {r^3} = \hfill \\ \hfill \\ = \frac{{12.56}}{3} \times {4^3} = \frac{{12.56}}{3} \times 64 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

2,411.25

OpenStudy (michele_laino):

I got this: \[\begin{gathered} V = \frac{{4 \times 3.14}}{3}{r^3} = \frac{{12.56}}{3} \times {r^3} = \hfill \\ \hfill \\ = \frac{{12.56}}{3} \times {4^3} = \frac{{12.56}}{3} \times 64 = 267.85c{m^3} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so 267.9

OpenStudy (michele_laino):

sorry I have made an error, here is the right result: \[\begin{gathered} V = \frac{{4 \times 3.14}}{3}{r^3} = \frac{{12.56}}{3} \times {r^3} = \hfill \\ \hfill \\ = \frac{{12.56}}{3} \times {4^3} = \frac{{12.56}}{3} \times 64 = 267.9466c{m^3} \hfill \\ \end{gathered} \] so the total volume of the requested chocolate is: \[V = 100 \times 267.9466 = 26794.66 \cong 26794.7c{m^3}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so rounded to the nearest foot it is 267.9

OpenStudy (michele_laino):

267.9 is the volume of a single volleyball

OpenStudy (michele_laino):

the requested volume of needed white chocolate is: \[V \cong 26794.7c{m^3}\]

OpenStudy (anonymous):

The new business building uptown is building a wall made entirely of glass pyramids. Each pyramid will have a square base with length of 8 inches and height of 10 inches. If the wall requires 250 of these pyramids, what volume of glass will be used? Round to the nearest cubic inch.

OpenStudy (michele_laino):

we have to compute the volume V of a single pyramid. So, we can apply this formula: \[\Large V = \frac{1}{3}L \times L \times H\] where L is the length of the square base, and H is the height of the pyramid: 

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

substituting your data, we have: \[\begin{gathered} V = \frac{1}{3}L \times L \times H = \hfill \\ \hfill \\ = \frac{1}{3}8 \times 8 \times 10 = \frac{{640}}{3} = ...inche{s^3} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

213.333333

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

Now, we have 250 of such pyramids, so the requested volume of needed glass, is: \[{V_{TOTAL}} = 213.3333 \times 250 = \]

OpenStudy (anonymous):

53,33.333333

OpenStudy (michele_laino):

that's right! Now we have to round off that result to the nearest unit, so we get: \[{V_{TOTAL}} = 213.3333 \times 250 = {\text{53333}}{\text{.325}} \cong {\text{53333}}inche{s^3}\]

OpenStudy (anonymous):

ok